19.5g CH3FCOOH is dissolved in 500g of water the depression in the freezing point observed is 1.0k calculate van't hoff factor and dissociation constant of the acid.
1
Expert's answer
2020-04-28T12:25:01-0400

Solution.

"\Delta T = i \times K(cr) \times Cm"

"Cm = \frac{19.5}{78.04 \times 0.5} = 0.50 \ mol/kg"

"i = \frac{\Delta T}{K(cr) \times Cm} = 1.08"

"\alpha = \frac{i-1}{k-1} = 0.08"

"Ka = \frac{\alpha^2C}{1- \alpha}"

"Ka = 3.48 \times 10^{-3}"

Answer:

i = 1.08

"Ka = 3.48 \times 10^{-3}"